# Invariants of persistence modules defined by order-embeddings

### v2

@@ 5.3. Comparing $\mathbf{mult}$-invariants
where $\beta^{\prime}_{(h,f,U)}=\alpha^{\prime}_{(h_{1},U)}$ as soon as $f=h_{2}$. ∎
@@
The following examples demonstrate how one reduces redundancy by choosing the set $\mathcal{F}$ given by increasingly larger posets. The first example addresses the problem raised in remark 3.9.
  1. (1)
    Let =$\mathcal{F}=\{(\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}})\}$ given as follows.
[…]
    Let $\mathcal{X}=\mathcal{X}_{2}$. Then we have two embeddings of $\mathcal{X}_{1}$ in $\mathcal{X}$ and one of $\mathcal{X}_{2}$ to $\mathcal{X}$. Moreover we have for $M$ in $\operatorname{\mathrm{mod}}\mathcal{X}$
    $\left(\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}(M)\right)_{f}=\left\{\begin{array}[]{cl}\dim M_{x}&\textrm{ if }f\in\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}_{2}}\textrm{ with }\operatorname{\mathrm{Im}}f=\{x\}\\ \operatorname{\mathrm{rk}}M_{1\to 2}&\textrm{ if }f\in\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}_{2}}\end{array}\right.$
    Then we immediately obtain
    $\mathbf{mult}_{\mathcal{X}}\sim\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}.$
    It is clear in this case that the composition map $\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{X}_{2}}\times\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ is surjective if $\mathcal{P}$ is connected, hence we obtain once again that
    $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}\simeq\mathbf{mult}_{\mathcal{F}}=\operatorname{\mathbf{rk}}_{\mathcal{P}}.$
  2. (2)
    Let $\mathcal{F}$ be the following family of posets :
[…]
    Set $\mathcal{X}=\mathcal{X}^{\prime}_{3}$.
    Then we have
    $\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}}=\{f_{1},f_{2},f_{3}\},\quad\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}}=\{f_{12},f_{13}\},\quad\textrm{ and }\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{X}}=\{{\rm id}\}.$
    Then one easily checks that we have
    $\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}=\phi\circ\mathbf{mult}_{\mathcal{X}}$
    with
    $\phi=\begin{pmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&1&1&0&0\\ 1&1&1&0&1&0\\ 1&1&1&1&1&1\end{pmatrix}$
    in the basis $\operatorname{\mathrm{Ind}}_{\mathcal{X}}=(\mathbb{I}_{\{1\}},\mathbb{I}_{\{2\}},\mathbb{I}_{\{3\}},\mathbb{I}_{[1,2]},\mathbb{I}_{[1,3]},\mathbb{I}_{\mathcal{X}^{\prime}_{3}}).$
  3. (3)
    Let $\mathcal{X}$ be the poset given by
    $a$$b$$c$$d$
    And let $\mathcal{F}$ be the family
[…]
    Then one can check that
    $|\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}}|=4,\ |\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}}|=5,\ |\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{X}}|=1,\ |\mathrm{Emb}_{\mathcal{X}^{\prime\prime}_{3}}^{\mathcal{X}}|=1.$
    The rank of the invariant $\mathbf{mult}_{\mathcal{X}}$ is also $11$ since there are $11$ indecomposable $\mathcal{X}$ modules. In a similar fashion, one can check that
    $\mathbf{mult}_{\mathcal{X}}\sim\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}.$
    Using Theorem 5.11 we obtain a basis of the invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$.
@@
    and let $\mathcal{X}:=\mathcal{X}_{4}.$ Then one can check that
@@
As a corollary of Example (2) we obtain the following
Theorem 5.15.
[…]
Let $\mathrm{E}=\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$ and recall that $\mathrm{I}=\operatorname{\mathrm{Ind}}\mathcal{X}$. Assume the following
[…]
Proof.
We consider the family of embeddings :=$\mathcal{F}:=\left\{(\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}),(\mathcal{X}_{3},\mathrm{Emb}_{\mathcal{X}_{3}}^{\mathcal{P}})\right\}$ where
[…]
and $\mathcal{X}_{3}=\mathcal{X}$. Because of condition ()(*)( ∗ ), the composition map $\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}}\times\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\to\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}$ is surjective. By condition ()(*)( ∗ ) and since $\mathcal{P}$ is connected, the composition map $\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}}\times\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\to\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}$ is also surjective. Then we can deduce from Example 2.6(2) that $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ is equivalent to $\mathbf{mult}_{\mathcal{F}}$. From Theorem 5.11, we deduce that there there is a basis of the form
$\{\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}}\mid f\in\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}},i=1,2,3\}.$
This basis contains all simples, all rectangles, and all generalized rectangles as described in Example 5.6 (2). One easily checks that the set $\mathcal{R}_{\mathrm{E}}$ contains exactly the simples, the rectangles and the generalized rectangles. The fact that it contains all the rectangles comes from condition ()(*)( ∗ ). ∎

### v3

@@ 5.3. Comparing $\mathbf{mult}$-invariants
where $\beta^{\prime}_{(h,f,U)}=\alpha^{\prime}_{(h_{1},U)}$ when $f=h_{2}$. ∎
We conclude by collecting several consequences of Proposition 5.14.
Proposition 5.15.
Let $\mathcal{P}$ be a finite connected poset and consider the following families of order embeddings
$\mathcal{F}\mathrel{\mathop{\ordinarycolon}}=((\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}),(\mathcal{X}^{\prime}_{3},\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}})),\quad\left(\textrm{resp.}\ \mathcal{F}\mathrel{\mathop{\ordinarycolon}}=((\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}),(\mathcal{X}^{\prime\prime}_{3},\mathrm{Emb}_{\mathcal{X}^{\prime\prime}_{3}}^{\mathcal{P}}))\right),$
where :
[…]
For $\mathcal{X}=\mathcal{X}^{\prime}_{3}$ (resp. $\mathcal{X}=\mathcal{X}^{\prime\prime}_{3}$) assume the following
[…]
Then we have an equivalence between the invariants $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ and $\mathbf{mult}_{\mathcal{F}^{\prime}}$.
Proof.
We prove the statement for $\mathcal{X}=\mathcal{X}^{\prime}_{3}$; the proof is similar for $\mathcal{X}=\mathcal{X}^{\prime\prime}_{3}$. We consider the family
$\mathcal{F}_{\mathcal{X}}=((\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}}),\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}}),\mathcal{X}^{\prime}_{3},\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{X}}))$
and show the equivalence between $\mathbf{mult}_{\mathcal{X}}$ and $\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}$.
We have
$\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}}=\{f_{1},f_{2},f_{3}\},\quad\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}}=\{f_{12},f_{13}\},\quad\textrm{ and }\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{X}}=\{{\rm id}\}$
where $f_{a}(1)=i$, $f_{ab}(1)=a$ and $f_{ab}(2)=b$. Then one easily checks that we have a commutative square
${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{X}}}$${\mathbb{Z}^{\{\mathbb{I}_{\{1\}},\mathbb{I}_{\{2\}},\mathbb{I}_{\{3\}},\mathbb{I}_{[1,2]},\mathbb{I}_{[1,3]},\mathbb{I}_{\mathcal{X}^{\prime}_{3}}\}}}$${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}}$${\mathbb{Z}^{\{f_{1},f_{2},f_{3},f_{12},f_{13},{\rm id}\}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$
with
$\Phi=\begin{pmatrix}1&0&0&-1&-1&1\\ 0&1&0&-1&0&0\\ 0&0&1&0&-1&0\\ 0&0&0&0&1&-1\\ 0&0&0&1&0&-1\\ 0&0&0&0&0&1\end{pmatrix}$.
Now condition ()(*) and the connectedness of $\mathcal{P}$ ensures that the composition map $\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{X}}\times\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\to\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}}$ is surjective for any $i$, and we can conclude using Proposition 5.14
Corollary 5.16.
$\mathcal{X}^{\prime}_{3}=$$\mathcal{X}^{\prime\prime}_{3}=112233$
Assume the following
[…]
Proof.
This is a direct consequence of the above proposition with Theorem 5.11. Note that the basis
$\{\mathbb{I}_{\overline{\operatorname{\mathrm{Im}}f}}\mid f\in\mathrm{Emb}_{\mathcal{X}_{i}}^{\mathcal{P}},i=1,2,3\}$
contains all simples, all rectangles, and all generalized rectangles as described in Example 5.6 (2). One easily checks that the set $\mathcal{R}_{\mathrm{E}}$ contains exactly the simples, the rectangles and the generalized rectangles. The fact that it contains all the rectangles comes from condition ()(*)
Theorem 5.17.
Let $\mathcal{P}$ be a finite connected poset, and let $\mathcal{X}$ be the following poset
[…]
and $\mathcal{X}_{1}$, $\mathcal{X}_{2}$, $\mathcal{X}^{\prime}_{3}$ and $\mathcal{X}^{\prime\prime}_{3}$ as above. Denote by $\mathrm{E}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Assume that the composition map
$\mathrm{Emb}_{\mathcal{Y}}^{\mathcal{X}}\times\mathrm{E}\longrightarrow\mathrm{Emb}_{\mathcal{\mathcal{Y}}}^{\mathcal{P}}$
is surjective for $\mathcal{Y}=\mathcal{X}_{1}$, $\mathcal{X}_{2}$, $\mathcal{X}^{\prime}_{3}$ and $\mathcal{X}^{\prime\prime}_{3}$. Then the set $\mathcal{R}_{\mathrm{E}}=\{\Theta_{f}(U)\mid(f,U)\in\mathrm{E}\times\operatorname{\mathrm{Ind}}(\mathcal{X})\}$ forms a basis of $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$.
Proof.
We introduce the family of embeddings
$\mathcal{F}=\left((\mathcal{X}_{1},\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{P}}),(\mathcal{X}_{2},\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{P}}),(\mathcal{X}^{\prime}_{3},\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{P}}),(\mathcal{X}^{\prime\prime}_{3},\mathrm{Emb}_{\mathcal{X}^{\prime\prime}_{3}}^{\mathcal{P}})\right)$
and
$\mathcal{F}_{\mathcal{X}}=\left((\mathcal{X}_{1},\mathrm{E}_{1}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Emb}_{\mathcal{X}_{1}}^{\mathcal{X}}),(\mathcal{X}_{2},\\ rE_{2}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Emb}_{\mathcal{X}_{2}}^{\mathcal{X}}),(\mathcal{X}^{\prime}_{3},\mathrm{E}^{\prime}_{3}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Emb}_{\mathcal{X}^{\prime}_{3}}^{\mathcal{X}}),(\mathcal{X}^{\prime\prime}_{3},\mathrm{E}^{\prime\prime}_{3}\mathrel{\mathop{\ordinarycolon}}=\mathrm{Emb}_{\mathcal{X}^{\prime\prime}_{3}}^{\mathcal{X}})\right).$
We first prove the equivalence between $\mathbf{mult}_{\mathcal{X}}$ and $\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}$. Indeed one can check that
$|\mathrm{E}_{1}|=4,\ |\mathrm{E}_{2}|=5,\ |\mathrm{E}^{\prime}_{3}|=1,\ |\mathrm{E}^{\prime\prime}_{3}|=1$
and construct an invertible map $\Phi\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{\mathrm{E}_{1}\sqcup\mathrm{E}_{2}\sqcup\mathrm{E}^{\prime}_{3}\sqcup\mathrm{E}^{\prime\prime}_{3}}$ to $\mathbb{Z}^{\operatorname{\mathrm{Ind}}\mathcal{X}}$ such that $\Phi\circ\mathbf{mult}_{\mathcal{F}_{\mathcal{X}}}=\mathbf{mult}_{\mathcal{X}}$. Then we can apply Proposition 5.14 to obtain an equivalence
$\mathbf{mult}_{\mathcal{F}}\sim\mathbf{mult}_{\mathcal{X},\mathcal{P}}.$
By Theorem 5.11 we obtain a basis of $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ by considering $\mathcal{P}$-modules of the form $\Theta_{f}(U)$ for $f\in\mathrm{E}_{1}$ (resp. $\mathrm{E}_{2}$, $\mathrm{E}^{\prime}3$, $\mathrm{E}^{\prime\prime}_{3}$) and $U=\mathbb{I}_{\mathcal{X}_{1}}$ (resp. $\mathbb{I}_{\mathcal{X}_{2}}$, $\mathbb{I}_{\mathcal{X}^{\prime}_{3}}$, $\mathbb{I}_{\mathcal{X}^{\prime\prime}_{3}}$). We can then check that these modules correspond to $\Theta_{f}(U)$ where $f\in\mathrm{E}$, and $U\in\operatorname{\mathrm{Ind}}\mathcal{X}$.
We finish by giving more examples of families of embeddings recovering known invariants using Proposition 5.14.
@@
    and let $\mathcal{X}\mathrel{\mathop{\ordinarycolon}}=\mathcal{X}_{4}.$ Then one can check that